# Can We Predict the Effect of Prompts?

### v1

@@ Abstract.
ccs: Software and its engineering Software creation and managementccs: Computing methodologies Machine learning
@@ 1. Introduction
Figure 1. An Overview of SPA, a Prototype Implementing Predictive Prompt Analysis
@@
As a demonstration that making such a prediction cheaply and without user-side training is possible, we showcase Syntactic Prevalence Analyzer (SPA), a preliminary predictive prompt analysis tool. We design SPA around a simple scenario: the user seeks to have the LLM generate a target syntactic structure, such as a try-except clause, during code synthesis, but there are multiple ways to ‘ask’ for such a structure; how effective would each prompt be in generating the target structure? SPA predicts an answer to this question, orders of magnitude more quickly than running the prompt on a code synthesis benchmark. SPA uses Sparse Autoencoders (SAEs), which are are models that ‘cluster’ the internal activation patterns of LLMs; specific clusters or ‘features’ of SAEs can often be mapped to recognizable concepts (Bricken et al., 2023). SPA first identifies the SAE features related to the user request, then based on these identified features predicts the relative incidence of the syntactic structures, ultimately allowing SPA to predict how well each prompt would meet the goal. Experiments demonstrate that SPA shows strong predictive performance – the actual incidence of the target syntactic structure for each prompt closely followed the predictions of SPA, with a Pearson correlation value of up to 0.994; meanwhile, the time it took to run the predictive analysis was only 0.4% of the total time to run the LLM on the full benchmark, and 18.7% of a baseline that generated 10 code samples per instruction, demonstrating the significant computational efficiency of our approach. Despite our strong early findings, there are two important limitations to SPA. First, in this preliminary work, the behavior we predict, syntactic structure prevalence, is fairly artificial. Second, our approach currently only works on open-weight LLMs for which there is a trained SAE. Nonetheless, we have optimistic initial results suggesting that improved predictive prompt analysis tools could overcome these limitations of SPA; such tools would ease rapid and computationally efficient prompt engineering. In summary, we (i) propose the concept of predictive prompt analysis; (ii) describe the prototype tool SPA implementing it; (iii) provide empirical results demonstrating the strong performance of SPA; and (iv) describe the future directions that are promising.
@@ 3. Methodology
Figure 2. Instructions asking for exception handling.
[…]
instructions
As described earlier, consider the scenario where a user wants a syntactic structure to be generated during LLM code synthesis, such that exception handlers are generated to prevent potential errors. The user is struggling between the options in Figure 2, as it is unclear a priori how consistently each instruction will lead to the LLM generating exception handling. Previously, the only way of answering this was to generate code based on these instructions hundreds or thousands of times. For example, even OpenAI suggests generating thousands of results to decide which prompt is better (OpenAI, nd), which is both computationally and financially expensive.
To mitigate this substantial cost, we propose SPA, a predictive prompt analysis technique which will predict which instruction is best. SPA takes two inputs: (1) the description of a target syntactic structure in natural language (e.g. “a try-except clause”) and (2) multiple prompts to compare, towards the goal of predicting which prompt will most consistently generate the target structure (shown in Fig. 2). With these inputs, SPA goes through two phases: (a) Extracting the Target Features and (b) Calculating the Ranking, as illustrated in Fig. 1, each described in detail next.
@@ 3.1. Extracting Target Features
To precisely extract the relevant SAE features, we employ ‘positive’ and ‘negative’ prompts. Figure 3 shows the template of a positive prompt, while removing “of <<OBJECTIVE>>” would make a negative prompt. Both prompts are needed to distinguish relevant SAE features related to the provided objective from generic features. For instance, on the prompt in Figure 3, SAE features that activate on "python code" or the word "write" are activated along with task-relevant features. Hence, a negative prompt, which has the same generic structure without mention of the objective, helps exclude generic features. We extract features that activate from the example code, highlighted in yellow, to find features that influence the LLM in response to the instruction.
@@ 4.2. Experimental Setup
Three syntactic objectives are used: generating try-except clauses, comments, and print statements. For each objective, both instruction sets and fewshot prompts were manually curated. We experiment with six prompt sets, also listed in Tab. 1. Exception and Print are composed of five instructions varying degrees of authoritative tone (see Fig. 2), while Comment is composed of eight instructions with three designed to suppress comment generation as the LLM tended to produce comments even when given no instructions. Exception Fewshot range from 0-shot to 4-shot, whereas Comment Fewshot and Print Fewshot are composed up to 3-shot due to memory constraints of our environment. The instructions for Exception can be found in Figure 2.
As a benchmark to evaluate our technique we used the sanitized version of Mostly Basic Python Problems (MBPP) which has 427 natural language specifications and corresponding code, a manually verified set of crowd-sourced Python programming problems (Austin et al., 2021). For SPA, we sampled 10 random problems ($|S|=10$) from MBPP and merged them with each instruction. As mentioned in Section 3.1, five target features were extracted from the prompts ($|TF|=5$). Pearson correlation was computed against LLM inference outputs on the entire MBPP dataset, averaged over three runs, with the total number of times the target syntactic structure was generated overall tallied. We also compare against a simple partial inference baseline that involves sampling 10 inference outputs from the same 10 problems used by SPA and counting the occurrence of the target syntactic structures. For both the partial inference and SPA predictions, the average was calculated on Fisher Z-transformed correlations over five attempts (Fisher, 1915).

### v2

# Predictive Prompt Analysis

@@ Abstract.
booktitle: Companion Proceedings of the 33rd ACM Symposium on the Foundations of Software Engineering (FSE ’25), June 23–27, 2025, Trondheim, Norwayconference: Companion Proceedings of the 33rd ACM Symposium on the Foundations of Software Engineering; June 23–27, 2025; Trondheim, Norwayccs: Software and its engineering Software creation and managementccs: Computing methodologies Machine learning
@@ 1. Introduction
Figure 1. An Overview of SPA, a Prototype Implementing Predictive Prompt Analysis. Upper Half (a): Extracting Target Features. Lower Half (b): Calculating the Ranking.
@@
As a demonstration that making such a prediction cheaply and without user-side training is possible, we showcase Syntactic Prevalence Analyzer (SPA), a preliminary predictive prompt analysis tool. We design SPA around a simple scenario: the user seeks to have the LLM generate a target syntactic structure, such as a try-except clause, during code synthesis, but there are multiple ways to ‘ask’ for such a structure; how effective would each prompt be in generating the target structure? SPA predicts an answer to this question, orders of magnitude more quickly than running the prompt on a code synthesis benchmark. SPA uses Sparse Autoencoders (SAEs), which are models that ‘cluster’ the internal activation patterns of LLMs; specific clusters or ‘features’ of SAEs can often be mapped to recognizable concepts (Bricken et al., 2023). SPA first identifies the SAE features related to the user request, then based on these identified features predicts the relative incidence of the syntactic structures, ultimately allowing SPA to predict how well each prompt would meet the goal. Experiments demonstrate that SPA shows strong predictive performance – the actual incidence of the target syntactic structure for each prompt closely followed the predictions of SPA, with a Pearson correlation value of up to 0.994; meanwhile, the time it took to run the predictive analysis was only 0.4% of the total time to run the LLM on the full benchmark, and 18.7% of a baseline that generated 10 code samples per instruction, demonstrating the significant computational efficiency of our approach. Despite our strong early findings, there are two important limitations to SPA. First, in this preliminary work, the behavior we predict, syntactic structure prevalence, is fairly artificial. Second, our approach currently only works on open-weight LLMs for which there is a trained SAE. Nonetheless, we have optimistic initial results suggesting that improved predictive prompt analysis tools could overcome these limitations of SPA; such tools would ease rapid and computationally efficient prompt engineering. In summary, we (i) propose the concept of predictive prompt analysis; (ii) describe the prototype tool SPA implementing it; (iii) provide empirical results demonstrating the strong performance of SPA; and (iv) describe the future directions that are promising.
@@ 3. Methodology
ID Instructions
1 None
2 It might be helpful to add an exception handler.
3 Write an exception handler.
4 You need to write an exception handler.
5 Please, with all my heart, include an exception handler.
Table 1. Exception Handler Instructions
As described earlier, consider the scenario where a user wants a syntactic structure to be generated during LLM code synthesis, such that exception handlers are generated to prevent potential errors. The user is struggling between the options in Table 1, as it is unclear a priori how consistently each instruction will lead to the LLM generating exception handling. Previously, the only way of answering this was to generate code based on these instructions hundreds or thousands of times. For example, even OpenAI suggests generating thousands of results to decide which prompt is better (OpenAI, nd), which is both computationally and financially expensive.
To mitigate this substantial cost, we propose SPA, a predictive prompt analysis technique which will predict which instruction is best. SPA takes two inputs: (1) the description of a target syntactic structure in natural language (e.g. “a try-except clause”) and (2) multiple prompts to compare, towards the goal of predicting which prompt will most consistently generate the target structure (shown in Tab. 1). With these inputs, SPA goes through two phases: (a) Extracting the Target Features and (b) Calculating the Ranking, as illustrated in Fig. 1, each described in detail next.
@@ 3.1. Extracting Target Features
To precisely extract the relevant SAE features, we employ ‘positive’ and ‘negative’ prompts. Figure 2 shows the template of a positive prompt, while removing “of <<OBJECTIVE>>” would make a negative prompt. Both prompts are needed to distinguish relevant SAE features related to the provided objective from generic features. For instance, on the prompt in Figure 2, generic SAE features that activate on "python code" or the word "write" are activated along with task-relevant features, limiting the effectiveness of the extraction process. Hence, a negative prompt, which has the same generic structure without mention of the objective, helps excluding these generic features. We extract features that activate from the example code, highlighted in yellow, to find features that influence the LLM in response to the instruction.
@@ 4.2. Experimental Setup
Three syntactic objectives are used: generating try-except clauses, comments, and print statements. For each objective, both instruction sets and fewshot prompts were manually curated through author consensus, simulating a prompt engineering process of involving instructions with varying tones. We experiment with six prompt sets, also listed in Tab. 2. Exception and Print are composed of five instructions varying degrees of authoritative tone (see Fig. 1), while Comment is composed of eight instructions with three designed to suppress comment generation as the LLM tended to produce comments even when given no instructions. For non-fewshot scenarios, an empty prompt was inserted to observe the inherent tendency of LLM generations. Exception Fewshot range from 0-shot to 4-shot, whereas Comment Fewshot and Print Fewshot are composed up to 3-shot due to memory constraints of our environment. The instructions for Exception can be found in Tab 1.
As a benchmark to evaluate our technique we used the sanitized version of Mostly Basic Python Problems (MBPP) which has 427 natural language specifications and corresponding code, a manually verified set of crowd-sourced Python programming problems (Austin et al., 2021). For SPA, we sampled 10 random problems ($|S|=10$) from MBPP and merged them with each instruction. As mentioned in Section 3.1, five target features were extracted from the prompts ($|TF|=5$). Pearson correlation for SPA’s predictions was computed against LLM inference outputs on the entire MBPP dataset, averaged over three runs, with the total number of times the target syntactic structure was generated overall tallied. A strong correlation would suggest that SPA effectively captures the causal dependencies between prompt and LLM behavior, thus validating its reliability as a quantitative metric for the task in question. We also compare against a partial inference baseline that involves sampling 10 inference outputs from the same 10 problems used by SPA and counting target syntactic structures. For both the partial inference and SPA predictions, the average was calculated on Fisher Z-transformed correlations over five attempts (Fisher, 1915).